Of 12 eggs in a refrigerator 2 are bad from these 4 eggs are choosen at random to make.what is the probility that at least one egg is bad?​
1
Expert's answer
2020-03-13T11:59:06-0400

Find the probability that the selected eggs are all good:

"P=\frac{C_{10}^4}{C_{12}^4}=\frac{14}{33}"

Then the probability that at least one is bad:

"\overline{P}=1-P=\frac{19}{33}"